would extend trade credit to some customers that previously have been denied credit because they were considered poor risks. Sales are projected to increase by $120,000 per year if credit is extended to these new customers. Of the new accounts receivable generated, 6% are projected to be uncollectible. Additional collection costs are projected to be 5% of incremental sales, and production and selling costs are projected to be 80% of sales. Your firm expects to pay a total of 30% of its income after expenses in taxes.
1) Compute the incremental income after taxes that would result from these projections:

Answer by Theo(13342)

additional gross income is 120,000.
6% is not collectible = .06 * 120,000 = 7,200.
5% is additional expense required for collecting = .05 * 120,000 = 6,000.
80% involves production and selling costs = .8 * 120,000 = 96,000.

total expenses are 7,200 + 6,000 + 96,000 = 109,200.
net income is therefore 120,000 - 109,200 = 10,800.
tax on that is 30% = 3,240.
incremental income is therefore 10,800 - 3,240 = 7,560.

that's what i get if all the costs are based on percentages of the incremental sales.